Badgersize Company has the following information for its Forming Department for the month of August Work in Process Inventory, August 1: 20,000 units Direct materials: 100% complete $80,000 24,000 $ 104,000 48,000 56,000 conversion: 20% complete Balance in work in process, August 1 Units started during August Units completed and transferred in August Work in process (70% complete), August 31 Costs charged to Work in Process in August Direct materials Conversion costs Direct labor Overhead applied $ 148,000 $ 117,000 146,000 $263,000 Total conversion Assume materials are added at the start of processing Required: a. Calculate the equivalent units for the Forming Department for the month of August. b. Find the cost per equivalent unit of input resource. (Round your answers to 2 decimal places.) Direct Materials Conversion a. Equivalent units b.Cost per equivalent unit

Explanations for each answer are appreciated.
Answer #1

Answer =

Direct material conversion
A equivalent units 48000 60400
B cost per equivalent units $3.08 $4.35

Explanation:

Units to account for physical units
Opening work in progress unit 20000
Unit started during the period 48000
Total unit to account for 68000
Units accounted for physical units Equivelent units (material) equivelent units (conversion cost)
Opening work in progress inventory 20000 16000(20000×80%)
Started and completed 36000 (56000-20000) 36000 36000
Closing work in progress inventory 12000 (20000 + 48000 - 56000) 12000 8400 (12000×70%)
Total units accounted for 48000 60400
Equivalent units 48000 60400

B)

Cost Per Equivalent Units (Material) = Material Cost/Equivalent Units for Material = 148000/48000 = $3.08

Cost Per Equivalent Units (Conversion) = Conversion Cost/Equivalent Units for Conversion = ($117000+146000)/60400 = $4.35

I have given complete and detailed solution of your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING